# Use of Nash equilibrium in finding game theoretic robust security bound on quantum bit error rate

### v1

@@ Acknowledgment:
We employ Wójcik’s attack strategy [66] as $E_{1}$ targeting the DL04 protocol in both message and control modes. As previously mentioned, Bob prepares the initial state randomly, choosing from $|0\rangle$, $|1\rangle$, |+ket|$\rangle$| + ⟩, and |ket|$\rangle$| - ⟩ with probabilities of $\frac{p}{2}$, $\frac{p}{2}$, $\frac{1-p}{2}$, and $\frac{1-p}{2}$, respectively. Alice encodes $j$ where $j\in\{0,1\}$ onto the travel photon [^8] Here, $t$ represents the travel photon (qubit) by Bob (Alice) for B-A attack (A-B attack). (indexed with the subscript $t$) received from Bob using operation [^9] Alice performs operation $iY_{t}^{0}\equiv I_{t}$ $\left(iY_{t}^{1}\equiv\left(ZX\right)_{t}^{1}\right)$ to encode 00 $\left(1\right)$ with probability $q$ $\left(1-q\right)$. $iY_{t}^{j}$. We assume that Alice and Bob utilize a lossy quantum channel with single-photon transmission efficiency denoted as $\eta$. In contrast, Eve replaces the channel with a perfect one with $\eta=1$. Eve employs two auxiliary spatial modes, $x$ and $y$, along with a single photon initially in the state $|0\rangle$. She conducts eavesdropping on the quantum channel twice: first during transmission from Bob to Alice (B-A attack) and then during transmission from Alice to Bob (A-B attack). The eavesdropping protocol (refer to Fig. 2 in Ref. [66]) begins by preparing two auxiliary modes, $x$ and $y$, in the state $|{\rm vac}\rangle_{x}|0\rangle_{y}$, where $|{\rm vac}\rangle$ represents an empty mode. Eve uses the unitary operation $Q_{txy}$ during the B-A attack and its conjugate, $Q_{txy}^{\dagger}\left(\equiv Q_{txy}^{-1}\right)$, during the A-B attack. The unitary operator $Q_{txy}(={\rm SWAP}_{tx}\,{\rm CPBS}_{txy}\,H_{y}\equiv{\rm SWAP}_{tx}\otimes I_{y}\,{\rm CPBS}_{txy}\,I_{t}\otimes I_{x}\otimes H_{y})$ is composed of a Hadamard gate (a one-qubit gate), a SWAP gate (a two-qubit gate), and a controlled polarizing beam splitter (CPBS), which is a three-qubit gate101010polarizing beam splitter which is assumed to transmit (reflect) photons in the state $|0\rangle$ $\left(|1\rangle\right)$.. We will now analyze this attack strategy for both message mode and control mode to derive the components of the payoff functions for Alice, Bob, and Eve.
Message mode: To begin with simplicity, let us consider the scenarios where Alice encodes 00. Specifically, we consider the situation where Bob prepares the initial state $|0\rangle_{t}$. We will then describe Eve’s attack on the composite system $|0\rangle_{t}|{\rm vac}\rangle_{x}|0\rangle_{y}$ is for $E_{1}$ attack strategy,
@@
This outcome aligns with the initial composite system, as anticipated because Alice employs an identity operation to encode 00, while Eve conducts a unitary A-B attack operation $\left(Q_{txy}\right)$ and its inverse, the unitary B-A attack operation $\left(Q_{txy}^{-1}\right)$. Likewise, we can describe the remaining composite systems in which Alice encodes 00 when Bob prepares initial states $|1\rangle_{t}$, $|+\rangle_{t}$, and $|-\rangle_{t}$, viz.,
@@
Eve’s optimal strategy Eve successfully decodes $k=0$ if her $x$ ancillary spatial mode is in the empty state and her $y$ ancillary spatial mode is in the state $|0\rangle$, which can be represented as $|{\rm vac}\rangle_{x}|0\rangle_{y}$. She decodes $k=1$ when she gets nonempty state in her $x$ ancillary spatial mode or empty state and $|1\rangle$ in her $x$ and $y$ ancillary spatial modes, respectively, i.e., $|0\rangle_{x}|{\rm vac}\rangle_{y}$ and $|1\rangle_{x}|{\rm vac}\rangle_{y}$ or $|{\rm vac}\rangle_{x}|1\rangle_{y}$. Now, let’s consider a parameter $p_{jmk}^{E_{1}}$, which characterizes the joint probability where $j$, $m$, and $k$ represent Alice, Bob, and Eve’s encoding, decoding, and decoding information, respectively. We can write the joint probabilities using Eqs. (6) -- (13) as follows.
@@
We can determine the mutual information among Alice, Bob, and Eve by applying Eq. (14) . To simplify the resulting expressions, we utilize Shannon entropy represented as, $\mathbf{H}\left[x\right]=-x\log_{2}x$, where $x$ represents the probability of an event occurring.
@@
Here also Alice encounters no detection of photon with probability $\frac{1}{2}$ like $Z$ basis scenarios but for all cases, the detection photon is not the same as Bob’s initial state (here $|+\rangle_{t}$). This implies that in B-A attack $\left(Q_{txy}\right)$, Eve’s detection happens for $X$ basis scenarios which do not occur in $Z$ basis cases. Now, we elaborately analyze Eve’s A-B attack $\left(Q_{txy}^{-1}\right)$ corresponding to all four cases of Alice’s collapse on measurement outcome. The final composite system on which Bob performs his measurement, given the collapsed system after Alice’s measurement is $|+\rangle_{t}|0\rangle_{x}|{\rm vac}\rangle_{y}$,
@@
Similarly, we can describe the remaining composite systems in which Alice encodes $1$ when Bob’s initial states are $|1\rangle$, |+ket|$\rangle$| + ⟩, and |ket|$\rangle$| - ⟩, i.e.,
@@
Eve’s optimal strategy The optimal strategy for Eve to decipher Alice’s encoded information, following her measurement of the $x$ and $y$ ancillary spatial modes, aligns with the $E_{1}$ attack scenario. We have already mentioned that with $\frac{1}{2}$ probability of the additional unitary operation $S_{ty}$ is performed. Consequently, Eq. (14) remains applicable for computing joint probabilities in half of all conceivable scenarios. These joint probabilities, denoted as $p_{jmk}^{E_{2}}$, where $j$, $m$, and $k$ represent Alice, Bob, and Eve’s encoding, decoding, and decoding information, respectively. We can write the joint probabilities using Eqs. (14) and (26) -- (33),
@@
By utilizing Equation (34), we can derive the mutual information expression among Alice, Bob, and Eve. This can be represented as follows,
@@
Eve first applies Hadamard gates to the $x$ and $y$ mode, then two CNOT gates, a polarizing beam splitter (PBS) gate, and two more CNOT gates (see Ref. [67] for more details).
@@
Henceforth, we will only provide the key results. Let’s consider the scenario where Bob prepares $|1\rangle_{t}$, $|+\rangle_{t}$, and $|-\rangle_{t}$. Following the B-A attack, the composite system becomes,
@@
The composite system after Eve’s attack operation $Q_{txy}^{\prime-1}$ (A-B attack) is as follows: When Alice encodes $1$ by applying the operation $iY_{t}^{1}$ to the Bob’s initial states $|1\rangle_{t}$, $|+\rangle_{t}$, and $|-\rangle_{t}$,
@@
The final composite systems resulting from Eve’s two-way attacks (B-A attack and A-B attack) when Alice encodes 00 (using the $I_{t}$ operation) with Bob’s initial states, $|0\rangle_{t}$, $|1\rangle_{t}$, $|+\rangle_{t}$, and $|-\rangle_{t}$ are,
@@
Eve’s optimal strategy Eve’s decoding process is characterized by two scenarios: one in which she decodes $k=0$ when her ancillary spatial modes $x$ and $y$ are in the empty state and $|0\rangle$ state, respectively, i.e., $|{\rm vac}\rangle_{x}|0\rangle_{y}$, and $k=1$ when she decodes with the $x$ mode in $|0\rangle$ state and $y$ mode in empty state, i.e., $|0\rangle_{x}|{\rm vac}\rangle_{y}$. These scenarios are described using a parameter $p_{jmk}^{E_{3}}$, which quantifies the joint probability, where $j$, $m$, and $k$ represent Alice, Bob, and Eve’s encoding, decoding, and decoding information, respectively. We can write the joint probabilities using Eqs. (46) -- (53),
@@
By utilizing Equation (54), we can derive the mutual information expression among Alice, Bob, and Eve. This can be represented as follows,
@@
Similarly, when Bob’s initial state is $|+\rangle_{t}$, and Alice measures in the $X$ basis, the composite system $|B-A\rangle_{|+\rangle E_{3}}$ collapses to one of the following states with equal probability: $|+\rangle_{t}|0\rangle_{x}|{\rm vac}\rangle_{y}$, $|-\rangle_{t}|0\rangle_{x}|{\rm vac}\rangle_{y}$, $|+\rangle_{t}|{\rm vac}\rangle_{x}|1\rangle_{y}$, $|-\rangle_{t}|{\rm vac}\rangle_{x}|1\rangle_{y}$, $|+\rangle_{t}|{\rm vac}\rangle_{x}|0\rangle_{y}$, $|-\rangle_{t}|{\rm vac}\rangle_{x}|0\rangle_{y}$, $|+\rangle_{t}|1\rangle_{x}|{\rm vac}\rangle_{y}$, and $|-\rangle_{t}|1\rangle_{x}|{\rm vac}\rangle_{y}$. Subsequently, the composite systems after Eve’s $Q_{txy}^{\prime-1}$ attack (A-B attack) are described,
@@
First, we define Eve’s IR attack ($E_{4}$ attack), afterward, we compute mutual information among Alice, Bob, and Eve. As B-A attack, Eve carries out a uniform (random) measurement with $Z$ and $X$ bases on Bob’s qubit and sends the projected qubit to Alice. either her encoding operation (in message mode) or her measurement operation (in control mode) on the received qubit and then sends it back to Bob. In the A-B attack, Eve applies the same measurement operation to the qubit that she had previously employed in the B-A attack.
Message mode: After performing a tedious calculation, we can obtain the joint probability, $p_{jmk}$, where $j$, $m$, and $k$ represent Alice, Bob, and Eve’s encoding, decoding, and decoding information, respectively.
@@
By utilizing Equation (63), we can derive the mutual information expression among Alice, Bob, and Eve. This can be represented as follows,
@@
$E_{1}$-$E_{2}$ subgame Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{1}-E_{2}}$
@@
$E_{1}$-$E_{3}$ subgame Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{1}-E_{3}}$
@@
$E_{2}$-$E_{3}$ subgame Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{2}-E_{3}}$
@@
$E_{1}$-$E_{4}$ subgame Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{1}-E_{4}}$
@@
Table 1: This table provides information on the Nash equilibrium points for various subgame scenarios and the qualitative payoff values for all parties.

### v2

@@ Acknowledgment:
We employ Wójcik’s attack strategy [73] as $E_{1}$ targeting the DL04 protocol in both message and control modes. As previously mentioned, Bob prepares the initial state randomly, choosing from $|0\rangle$, $|1\rangle$, |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ with probabilities of $\frac{p}{2}$, $\frac{p}{2}$, $\frac{1-p}{2}$ and $\frac{1-p}{2}$, respectively. Alice encodes $j$ where $j\in\{0,1\}$ onto the travel photon [^10] Here, $t$ represents the travel photon (qubit) by Bob (Alice) for B-A attack (A-B attack). (indexed with the subscript $t$) received from Bob using operation [^11] Alice performs operation $iY_{t}^{0}\equiv I_{t}$ $\left(iY_{t}^{1}\equiv\left(ZX\right)_{t}^{1}\right)$ to encode 00 $\left(1\right)$ with probability $q$ $\left(1-q\right)$. $iY_{t}^{j}$. We assume that Alice and Bob utilize a lossy quantum channel with single-photon transmission efficiency denoted as $\eta$. In contrast, Eve replaces the channel with a perfect one with $\eta=1$. Eve employs two auxiliary spatial modes, $x$ and $y$, along with a single photon initially in the state $|0\rangle$. She conducts eavesdropping on the quantum channel twice: first during transmission from Bob to Alice (B-A attack) and then during transmission from Alice to Bob (A-B attack). The eavesdropping protocol (refer to Fig. 2 in Ref. [73]) begins by preparing two auxiliary modes, $x$ and $y$, in the state $|{\rm vac}\rangle_{x}|0\rangle_{y}$, where $|{\rm vac}\rangle$ represents an empty mode. Eve uses the unitary operation $Q_{txy}$ during the B-A attack and its conjugate, $Q_{txy}^{\dagger}\left(\equiv Q_{txy}^{-1}\right)$, during the A-B attack. The unitary operator $Q_{txy}(={\rm SWAP}_{tx}\,{\rm CPBS}_{txy}\,H_{y}\equiv{\rm SWAP}_{tx}\otimes I_{y}\,{\rm CPBS}_{txy}\,I_{t}\otimes I_{x}\otimes H_{y})$ is composed of a Hadamard gate (a one-qubit gate), a SWAP gate (a two-qubit gate) and a controlled polarizing beam splitter (CPBS), which is a three-qubit gate121212polarizing beam splitter which is assumed to transmit (reflect) photons in the state $|0\rangle$ $\left(|1\rangle\right)$.. We will now analyze this attack strategy for both message mode and control mode to derive the components of the payoff functions for Alice, Bob and Eve.
Message mode: To begin with simplicity, let us focus on scenarios where Alice encodes 0. Specifically, we will examine the situation where Bob prepares the initial state $|0\rangle_{t}$. We will then discuss Eve’s attack strategy on the composite system, which involves the state $|0\rangle_{t}|{\rm vac}\rangle_{x}|0\rangle_{y}$ and is denoted as $E_{1}$,
@@
This outcome aligns with the initial composite system, as anticipated because Alice employs an identity operation to encode 00, while Eve conducts a unitary A-B attack operation $\left(Q_{txy}\right)$ and its inverse, the unitary B-A attack operation $\left(Q_{txy}^{-1}\right)$. Likewise, we can describe the remaining composite systems in which Alice encodes 00 when Bob prepares initial states $|1\rangle_{t}$, $|+\rangle_{t}$ and $|-\rangle_{t}$, viz.,
@@
Eve’s optimal strategy Eve successfully decodes $k=0$ if her $x$ ancillary spatial mode is in the empty state and her $y$ ancillary spatial mode is in the state $|0\rangle$, which can be represented as $|{\rm vac}\rangle_{x}|0\rangle_{y}$. She decodes $k=1$ when she gets nonempty state in her $x$ ancillary spatial mode or empty state and $|1\rangle$ in her $x$ and $y$ ancillary spatial modes, respectively, i.e., $|0\rangle_{x}|{\rm vac}\rangle_{y}$ and $|1\rangle_{x}|{\rm vac}\rangle_{y}$ or $|{\rm vac}\rangle_{x}|1\rangle_{y}$. Now, let’s consider a parameter $p_{jmk}^{E_{1}}$, which characterizes the joint probability where $j$, $m$ and $k$ represent Alice, Bob and Eve’s encoding, decoding, and decoding information, respectively. We can write the joint probabilities using Eqs. (6) -- (13) as follows.
@@
We can determine the mutual information among Alice, Bob and Eve by applying Eq. (14) . To simplify the resulting expressions, we utilize Shannon entropy represented as, $\mathbf{H}\left[x\right]=-x\log_{2}x$, where $x$ represents the probability of an event occurring.
@@
Alice also encounters no detection of a photon with a probability of $\frac{1}{2}$, similar to the $Z$ basis scenarios. However, in all detected cases, the photon is not in Bob’s initial state ($|+\rangle_{t}$). This implies that in the B-A attack $\left(Q_{txy}\right)$, Eve’s detection occurs for $X$ basis scenarios, which does not happen in $Z$ basis cases. Now, we will thoroughly analyze Eve’s A-B attack $\left(Q_{txy}^{-1}\right)$ for all four possible outcomes of Alice’s measurement outcome. The final composite system on which Bob performs his measurement, given the collapsed state after Alice’s measurement, is $|+\rangle_{t}|0\rangle_{x}|{\rm vac}\rangle_{y}$,
@@
Similarly, we can describe the remaining composite systems in which Alice encodes $1$ when Bob’s initial states are $|1\rangle$, |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩, i.e.,
@@
Eve’s optimal strategy The optimal strategy for Eve to decipher Alice’s encoded information, following her measurement of the $x$ and $y$ ancillary spatial modes, aligns with the $E_{1}$ attack scenario. We have already mentioned that with $\frac{1}{2}$ probability of the additional unitary operation $S_{ty}$ is performed. Consequently, Eq. (14) remains applicable for computing joint probabilities in half of all conceivable scenarios. These joint probabilities, denoted as $p_{jmk}^{E_{2}}$, where $j$, $m$ and $k$ represent Alice, Bob and Eve’s encoding, decoding and decoding information, respectively. We can write the joint probabilities using Eqs. (14) and (26) -- (33),
@@
By utilizing Equation (34), we can derive the mutual information expression among Alice, Bob and Eve. This can be represented as follows,
@@
Eve first applies Hadamard gates to the $x$ and $y$ mode, then two CNOT gates, a polarizing beam splitter (PBS) gate and two more CNOT gates (see Ref. [74] for more details).
@@
Henceforth, we will only provide the key results. Let’s consider the scenario where Bob prepares $|1\rangle_{t}$, $|+\rangle_{t}$ and $|-\rangle_{t}$. Following the B-A attack, the composite system becomes,
@@
The composite system after Eve’s attack operation $Q_{txy}^{\prime-1}$ (A-B attack) is as follows: When Alice encodes $1$ by applying the operation $iY_{t}^{1}$ to the Bob’s initial states $|1\rangle_{t}$, $|+\rangle_{t}$ and $|-\rangle_{t}$,
@@
The final composite systems resulting from Eve’s two-way attacks (B-A attack and A-B attack) when Alice encodes 00 (using the $I_{t}$ operation) with Bob’s initial states, $|0\rangle_{t}$, $|1\rangle_{t}$, $|+\rangle_{t}$ and $|-\rangle_{t}$ are,
@@
Eve’s optimal strategy Eve’s decoding process is characterized by two scenarios: one in which she decodes $k=0$ when her ancillary spatial modes $x$ and $y$ are in the empty state and $|0\rangle$ state, respectively, i.e., $|{\rm vac}\rangle_{x}|0\rangle_{y}$, and $k=1$ when she decodes with the $x$ mode in $|0\rangle$ state and $y$ mode in empty state, i.e., $|0\rangle_{x}|{\rm vac}\rangle_{y}$. These scenarios are described using a parameter $p_{jmk}^{E_{3}}$, which quantifies the joint probability, where $j$, $m$ and $k$ represent Alice, Bob and Eve’s encoding, decoding and decoding information, respectively. We can write the joint probabilities using Eqs. (46) -- (53),
@@
By utilizing Equation (54), we can derive the mutual information expression among Alice, Bob and Eve. This can be represented as follows,
@@
Similarly, when Bob’s initial state is $|+\rangle_{t}$, and Alice measures in the $X$ basis, the composite system $|B-A\rangle_{|+\rangle E_{3}}$ collapses to one of the following states with equal probability: $|+\rangle_{t}|0\rangle_{x}|{\rm vac}\rangle_{y}$, $|-\rangle_{t}|0\rangle_{x}|{\rm vac}\rangle_{y}$, $|+\rangle_{t}|{\rm vac}\rangle_{x}|1\rangle_{y}$, $|-\rangle_{t}|{\rm vac}\rangle_{x}|1\rangle_{y}$, $|+\rangle_{t}|{\rm vac}\rangle_{x}|0\rangle_{y}$, $|-\rangle_{t}|{\rm vac}\rangle_{x}|0\rangle_{y}$, $|+\rangle_{t}|1\rangle_{x}|{\rm vac}\rangle_{y}$ and $|-\rangle_{t}|1\rangle_{x}|{\rm vac}\rangle_{y}$. Subsequently, the composite systems after Eve’s $Q_{txy}^{\prime-1}$ attack (A-B attack) are described,
@@
First, we define Eve’s IR attack ($E_{4}$ attack), afterward, we compute mutual information among Alice, Bob and Eve. As B-A attack, Eve carries out a uniform (random) measurement with $Z$ and $X$ bases on Bob’s qubit and sends the projected qubit to Alice. either her encoding operation (in message mode) or her measurement operation (in control mode) on the received qubit and then sends it back to Bob. In the A-B attack, Eve applies the same measurement operation to the qubit that she had previously employed in the B-A attack.
Message mode: After performing a tedious calculation, we can obtain the joint probability, $p_{jmk}$, where $j$, $m$ and $k$ represent Alice, Bob and Eve’s encoding, decoding and decoding information, respectively.
@@
By utilizing Equation (63), we can derive the mutual information expression among Alice, Bob and Eve. This can be represented as follows,
@@
$E_{1}$-$E_{2}$ game scenario Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{1}-E_{2}}$
@@
$E_{1}$-$E_{3}$ game scenario Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{1}-E_{3}}$
@@
$E_{2}$-$E_{3}$ game scenario Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{2}-E_{3}}$
@@
$E_{1}$-$E_{4}$ game scenario Nash equilibrium point ($p,q,r$) Alice’s/Bob’s payoff Eve’s payoff Payoff difference $\epsilon_{E_{1}-E_{4}}$
@@
Table 2: This table provides information on the Nash equilibrium points for various game scenarios and the qualitative payoff values for all parties.